OpenStudy (anonymous):

can someone explain to me parametric equations(calc2)

OpenStudy (anonymous):

can you post afew examples?

OpenStudy (anonymous):

given the curve x=t^2-t+2, y=t^3-3t find the points of horizontal and vertical tangency

OpenStudy (anonymous):

think of t in seconds so equation 1 x=t^2-t+2 at t=5 x=25-5+2 x=18 with equation 2 you are y=t^3-3t y=125-15 =110  im sure you get asked to change the two equations into 1